EarthWear Clothiers Investor Website 1. Who is the current CEO of EarthWear? 2. How much did Accounts Payable change in dollars and percent from 2014 to 2015? 3 In which of the last 5 years was EarthWears Net Income highest, and what was Net Income in that year? 4. Does EarthWear use the FIFO or LIFO inventory valuation method? Hint See the Notes to the Consolidated Financial Statements Willis & Adams, CPAs Website 5. What 5 services do Willis & Adams, CPAs provide to clients? What are the names of the three current staff accountants listed on the Willis & Adams website? 7. What is the date of Willis & Adams most current engagement letter wit h EarthWear?
Answer #1
  1. The current CEO of EarthWear is Calvin J. Rodgers.
  2. Change in Accounts Payable = 62,509 – 48,432 = $14077
    % Change = 14077/62509 = 22.52%
  3. Net income was the highest in 2012 at $31,222.
  4. LIFO inventory valuation method
  5. Assurance Services, Tax Planning, Business Planning, Valuation Services, and Litigation Support Services.
  6. Michael Willis, Scott Adams, Karen Mitchell
  7. The most current engagement letter with EarthWear is February 15, 2016